OpenStudy (anonymous):

Help me. It's 3a.m. in the morning and I'm stuck on my last question. Prove: cos^2x-sin^2x/cosx+sinx = cscx - secx/secx cscx

OpenStudy (mathstudent55):

\( \dfrac{\cos^2x-\sin^2x}{\cos x+\sin x} = \dfrac{\csc x - \sec x}{\sec x \csc x} \) Is this it?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

I've tried both sides but it doesn't seem to work out...

OpenStudy (campbell_st):

well why not factor the numerator \[\frac{(\cos(x) + \sin(x))(\cos(x) - \sin(x))}{\sin(x) + \cos(x)}\]

OpenStudy (anonymous):

doesn't that just leave cosx-sinx

OpenStudy (campbell_st):

well use the fact that the numerator is cos(2x) and cos(2x) = 1 - 2sin^2(x) or cos(2x) - 2cos^2(x) - 1 or cos(2x) = (1 - tan^2(x))/(1 + tan^2(x))

OpenStudy (mathstudent55):

\(\large \dfrac{\cos^2x-\sin^2x}{\cos x+\sin x} = \dfrac{\csc x - \sec x}{\sec x \csc x}\) \(\large = \dfrac{\frac{1}{\sin x} -\frac{1}{\cos x} } {\frac{1}{\cos x}{\frac{1}{\sin x} }} \) \(\large = \dfrac{\frac{1}{\sin x} -\frac{1}{\cos x} } {\frac{1}{\cos x}{\frac{1}{\sin x} }} \times \dfrac{\sin x \cos x}{\sin x \cos x} \) \(\large = \dfrac{\cos x -\sin x} {1} \) \(\large = \dfrac{\cos x -\sin x} {1} \times \dfrac{\cos x + \sin x}{\cos x + \sin x} \) \(\large = \dfrac{\cos^2 x -\sin^2 x} {\cos x + \sin x} \)

OpenStudy (anonymous):

I have another question, is the problem: (secx+cscx)cotx = false? because everyone cant seem to solve it...

OpenStudy (mathstudent55):

You can certainly multiply that out. Is it supposed to be equal to something?

OpenStudy (campbell_st):

so you have 1/cos * cos/sin + 1/sin * cos/sin = 1/sin + cos/sin^2 so csc + csc*tan which becomes csc(1 + tan)

OpenStudy (anonymous):

oh it's supposed to be equal to 1

zepdrix (zepdrix):

Hmm these are all good methods :) This is how I would have done it though,\[\Large\bf\sf \frac{(\cos x-\sin x)(\cos x+\sin x)}{\cos x+ \sin x}\quad=\quad \frac{\csc x - \sec x}{\sec x \csc x}\] \[\Large\bf\sf \cos x - \sin x quad=\quad \frac{\csc x - \sec x}{\sec x \csc x}\] \[\Large\bf\sf \frac{1}{\sec x}-\frac{1}{\csc x} \quad=\quad \frac{\csc x - \sec x}{\sec x \csc x}\]Then just get a common denominator in the left side.

zepdrix (zepdrix):

Grr stupid latex fail t.t

OpenStudy (mathstudent55):

\( (\sec x + \csc x) \cot x~~~~~~~~~~~ = 1 \) \(\left( \dfrac{1}{\cos x} + \dfrac{1}{\sin x} \right) \dfrac{\cos x}{\sin x}~~~~ = 1\) \( \dfrac{1}{\sin x} + \dfrac{\cos x}{\sin^2 x} ~~~~~~~~~~~~~~~~~=1\) \( \dfrac{1}{\sin x} \times \dfrac{\sin x}{\sin x} + \dfrac{\cos x}{\sin^2 x} ~~~~~=1\) \(\dfrac{\sin x}{\sin^2 x} + \dfrac{\cos x}{\sin^2 x} ~~~~~~~~~~~~~~~~=1\) \(\dfrac{\sin x + \cos x}{\sin^2 x} ~~~~~~~~~~~~~~~~~~~~=1\) \(\sin x + \cos x = {\sin^2 x} \) It doesn't look like an identity to me. This may be true for certain values of x which makes this an equation, not an identity.

OpenStudy (anonymous):

Yea, it's okay, I knew it wouldn't work...But thanks a ton. I wish I could hug you and throw you some food but I can't, sadly... >n<

OpenStudy (mathstudent55):

LOL. You're welcome.